What is x if a=x, b=x+6, and c= 2 square roots of 17?
Answer by stanbon(75887)

What is x if a=x, b=x+6, and c= 2 square roots of 17?
-----
c^2 = a^2 + b^2
(2sqrt(17))^2 = x^2 + (x+6)^2
---------------------
4*17 = x^2 + x^2+12x+36
-------
2x^2 + 12x -32 = 0
-------
x^2 + 6x - 16 = 0
-------
(x+8)(x-2) = 0
--------
x = -8 or x = 2
--------------------
Cheers,
Stan H.
--------------